OpenStudy (anonymous):

How do you find the x-intercept of y=x^2+2x-3? I know you set it to 0 but I am not sure what to do after that

OpenStudy (zepp):

You either factor it out or use the quadratic formula ;)

OpenStudy (anonymous):

oh ok thank you :)

OpenStudy (lgbasallote):

"set it to zero" you do mean set y to zero right?

OpenStudy (anonymous):

you can factor it out

OpenStudy (anonymous):

yes thats what i meant

OpenStudy (anonymous):

you quadratic formula when u can factor out the numebers

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

cant*

OpenStudy (anonymous):

so if y = -3, and x = 1, how do you find the vertex? or whats the formula to find it

OpenStudy (zepp):

To find the vertex, use a point and zero Then plug in it \(\large y=a(x-z_{1})(x-z_{2})\)

OpenStudy (zepp):

Uh, forget what I said. :s

OpenStudy (anonymous):

@zepp lol ok i looked at it and was like uh...

OpenStudy (zepp):

But what's the (1,-3)?

OpenStudy (zepp):

A random point on the parabola?

OpenStudy (callisto):

Make it into the form: y=a(x-h)+k, where (h,k) are the coordinates of the vertex

OpenStudy (anonymous):

that was the first part of the problem was to find the y intercept which turned to be -3, and x intercept was 1 @zepp

OpenStudy (zepp):

A quadratic function must have 2 x-intercept or none.

OpenStudy (zepp):

There's 1 missing :)

OpenStudy (anonymous):

than the other is -3

OpenStudy (anonymous):

those were the two that i got

OpenStudy (callisto):

I suppose it's called completing of square... y=x^2+2x-3 = (x^2 + 2x +1) -1 -3 = (x+1)^2 - 4 Coordinates of vertex are (-1, -4)

OpenStudy (zepp):

..Alright, the equation was given. Thought it wasn't >.>